Copper forms two compounds with chlorine.

Compound A: 32.10 g Cu and 17.90 g Cl

Compound B: 23.64 g Cu and 26.37 g Cl






1
Expert's answer
2021-01-05T05:59:54-0500

Solution:

Molar mass of Cu 63.55 g/mol

Molar mass of CI 35.45 g/mol

Compound A

n Cu = 32.10g/63.55g/mol = 0.505 mol

n CI = 17.90g/35.45g/mol = 0.505 mol

Since atoms combine in the same ratio that moles do, we divide all of the numbers of moles by the smallest number to put everything into lowest terms:

Cu 0.505 Cl 0.505 → Cu 0.505/0.505 Cl 0.505/0.505 → CuCl

Compound B

n Cu = 23.64g/63.55g/mol = 0.372 mol

n CI = 26.37g/35.45g/mol = 0.744 mol

Cu 0.372 Cl 0.744 → Cu 0.372/0.372 Cl 0.744/0.372 → CuCl2

Answer: compound A empirical formula CuCl; compound B empirical formula CuCl2